OpenStudy (anonymous):

need to find the answer for x please solve and explain 0.9x+2.3x=-6.4

OpenStudy (anonymous):

1. Collect like terms (add 0.9x and 2.3x together!) 2. Divide each side by the coefficient of x

OpenStudy (anonymous):

this question is not for that its for finding what "x" equals to

OpenStudy (anonymous):

The procedure I gave you will give you what x equals to. 0.9x+2.3x=-6.4 3.1x=-6.4 x=-6.4/3.1

OpenStudy (anonymous):

where did you get the 3.1 from?

OpenStudy (anonymous):

0.9+2.3 = 3.1

OpenStudy (anonymous):

um doesnt it equal to 3.2....

OpenStudy (anonymous):

oops yes my bad :P

OpenStudy (anonymous):

ok so 0.9x+2.3x=-6.4 3.2x=-6.4 x=-6.4/3.2 x=2 ? correct

OpenStudy (anonymous):

Very close. -2

OpenStudy (anonymous):

oh yes sorry didnt notice the sign there

OpenStudy (anonymous):

thank you